OpenStudy (anonymous):

Please help! Given f(x)=xe^x, find f'(x), f''(x), f''(0), f'''(x) and f'''(0).

OpenStudy (anonymous):

f'(x)=(e^x)(x+1), f"(x)=(e^x)(x+2), f"(0)=2, f'''(x)=(e^x)(x+3), f'''(0)=3